[Dsg-hallb_magnets] Solenoid tests and modifications
Brian Eng
beng at jlab.org
Sat Sep 8 12:20:20 EDT 2018
The issue in 2.5 is that because of the bad opto the reset line was stuck getting pulled to always reset. Aside from that in and of itself being bad, the reset is common to all QD boards on a magnet, so for the solenoid both were stuck. Replacing that component should have fixed it (I don't know the extent of the repairs).
Onish said they skipped testing the units in the lab after the pot replacement since they would consider the tuning in-situ the testing.
This saves time since there was a bit of a time crunch and if the units are good no issues. Obviously it has a major flaw when there are problems with the boards.
We wouldn't have done any changing of the QDs if it weren't for the extended ACC downtime though.
New niece is Stella Mei, both are fine now. Apparently it's new procedure to keep them there for at least 2 days afterwards, not sure if it's hospital specific or everywhere.
----- Original Message -----
> From: "Amrit Yegneswaran" <yeg at jlab.org>
> To: "Pablo Campero" <campero at jlab.org>
> Cc: "dsg-hallb magnets" <dsg-hallb_magnets at jlab.org>
> Sent: Friday, September 7, 2018 9:58:41 PM
> Subject: Re: [Dsg-hallb_magnets] Solenoid tests and modifications
> pablo,
> thanks for the detailed write up.
> in point 2.5 of your report, what issues were found with the qd and how were
> these issues fixed.
>
> my question would be why were there any issues at all?
> were the spare qds not checked and tested ahead of time?
> it is indeed lucky that we were replacing the qds when there was no scheduled
> physics run in progress.
> hmm, indeed, indeed!
>
> amanda, brian, tyler, and you did a great job supporting probir and ruben.
> i'm sure they appreciate your support.
>
> i guess congratulations to the young uncle is in order, so what's her name
> brian?
> trust mother and child are doing well.
>
> wish you guys a great weekend.
> goodnight and godbless.
> amrit
>
>
>
> From: "Pablo Campero" <campero at jlab.org>
> To: "dsg-hallb magnets" <dsg-hallb_magnets at jlab.org>
> Sent: Friday, September 7, 2018 6:11:05 PM
> Subject: [Dsg-hallb_magnets] Solenoid tests and modifications
>
>
>
> Hello,
>
>
>
>
> To summarize the test and modifications performed on the Solenoid we can
> mention:
>
>
>
>
> 1. On 09/05/2018 it was performed a test to verify the timing between the PLC
> SOE and FastDAQ data/ FastDump switch opening.
>
> Solenoid ramped up to 100 A (two times) and fast dumped intentionally later to
> monitor the sequence of the events and the timestamps for the FastDAQ data.
> During this test QD#1 tripped first after the Fast_Dump button was pushed.
>
> As the outstanding of this test, it was found the following timing differences
> based in the SOE timestamps and FastDAQ data timestamps:
>
> Test #1:
>
> FastDAQ_QD1 - SOE _Fast Dump Relay was ~ 270.8 ms.
>
> FastDAQ_IDCCT1 - FastDAQ_QD1 was ~36.6 ms
>
> SOE_QD1 - FastDAQ_QD1 was ~ 34 ms.
>
> Test #2 :
>
> FastDAQ_QD1 - SOE _Fast Dump Relay was ~ 284.2 ms.
>
> FastDAQ_IDCCT1 - FastDAQ_QD1 was ~26.3 ms
>
> SOE_QD1 - FastDAQ_QD1 was ~ 34.7 ms.
>
>
>
> For more details about this tests please refer:
> https://logbooks.jlab.org/entry/3593687
>
>
>
>
> 2. Today 09/07/2018 it was performed the swapping of QD#1 and QD#2 units by two
> QD spares modified.
>
>
> Two spare Quench Detector units were modified by DC group with the aim to find
> more stability during the magnet operations. Modifications were applied to
> their hardware components as follow:
>
> - Changed eight potentiometers for each channel (2 potentiometers per channel,
> one is used for gain and the other for balance). Pots changed from 12 turns to
> 25 turns.
>
> - Modified LED voltage source wiring. LEDs used as indicators for QD-Ch
> lower/upper and Sum.
>
> The following steps were followed for the swapping of the QDs:
>
> 2.1. Ramped down the Solenoid to 0 A.
>
> 2.2. Connected a external power source to enable control power in the MPS.
>
> 2.3. Took off and disconnected QD#1 and QD#2 units from the controls rack.
>
> 2.4. Installed modified QDs, re-connected power cable, and signal/control
> cables.
>
> 2.5. Found issues with the modified QD replaced as QD#2. (fixed later by
> Onish/DC group- changed Octo/component on QD#2 board)
>
>
> 2.6. Tuned QD#1 and QD#2. Set up gain/ balance pots and set voltage thresholds
> as previous ones (0.1 and 1.5 V)
>
> 2.7. Recorded values at which QD channels tripped.
>
> 2.8. Set up delay time thresholds (200 and 999 ms)for each channel and make sure
> that all were recorded in the QDs memory.
>
> 2.9 Tested QDs by tripping them when thresholds were exceeded.
>
> 2.10. Removed control power for MPS and put it back to normal.
>
>
>
>
> After tunning was completed and returned all controls/instrumentation(VT panel)
> to normal operations the solenoid was ramped up to 250 A, it will be kept at
> this current over this weekend.
>
> The next step will be ramp the Solenoid up to full current (2416 A) on Monday
> and test the modifications made in QDs.
>
>
>
>
> Best Regard,
>
> Pablo
>
>
>
>
>
>
>
>
>
>
>
>
>
>
> _______________________________________________
> Dsg-hallb_magnets mailing list
> Dsg-hallb_magnets at jlab.org
> https://mailman.jlab.org/mailman/listinfo/dsg-hallb_magnets
>
> _______________________________________________
> Dsg-hallb_magnets mailing list
> Dsg-hallb_magnets at jlab.org
> https://mailman.jlab.org/mailman/listinfo/dsg-hallb_magnets
More information about the Dsg-hallb_magnets
mailing list